OpenStudy (anonymous):

Solve for both exact form and decimal form (rounded to 3 decimal places). :) 1= 14tan (6x+8) - 1 not sure... but do I do this? 1= 14tan(6x) + 14tan(8) -1 add 1 to both sides 2=14tan(6x) + 14tan(8) ?? :/ is that the correct way to start? :/ thankyouu:)

OpenStudy (anonymous):

no :D 1= 14tan (6x+8) - 1 /+1 2=14tan(6x+8) /14 1/7=tan(6x+8) Let t=6x+8 1/7=tan(t), solve simple equation per t and then you are solving linear equation t=somethin+k*pi something+k*pi=6x+8 -8+(something+k*pi)/6=x

OpenStudy (phi):

tan( stuff) are complicated

OpenStudy (anonymous):

ohh okay... i'm not sure what you mean by the t=something + k*pi part though.. i know the k is the constant right? :/ but how do we figure out the something ?

OpenStudy (anonymous):

1/7=tan(6x+8) Let t=6x+8 1/7=tan(t) t=arctan(1/7) (type on calculator -> t=0.1419+kPi, k is integer 0.1419+kPi=6x+8 -8+(0.1419+kPi)/6=x you can simplify this if you want to :)

OpenStudy (phi):

the tangent repeats every pi radians (every 180 deg) once you find an answer, you can add pi to it to get another answer (or subtract pi) in fact you can add any multiple of pi to get another answer.

OpenStudy (agent0smith):

1= 14tan (6x+8) - 1 1= 14tan(6x) + 14tan(8) -1 can't do this. Tan(a + b) does NOT equal tana + tanb. Same applies to sin and cos

OpenStudy (anonymous):

@purger -7.97635 + kπ/6 = x ? not sure if that's right haha :P and ohh okay @agent0smith oops :P

OpenStudy (anonymous):

@iheartfood 6x+8=0.14189+kpi 6x=-7.85...+kPi x=-1.309683+kPi/6, k is integer

OpenStudy (anonymous):

ohh okay.. so does that simplify any further? or is that what would be considered the exact form? :/ or is that the decimal form (since it's a reallyyyy long number haha) ?

OpenStudy (agent0smith):

If you needed exact form, which is odd in this case, just leave it as 6x+8 = tan^-1(1/7) and solve from there, same as above..

OpenStudy (anonymous):

so what id be solving for would result in the decimal form? :/

OpenStudy (agent0smith):

6x+8 = tan^-1(1/7) solve for x: x = (tan^-1(1/7) - 8)/6

OpenStudy (agent0smith):

No that's exact form, left as that.

OpenStudy (anonymous):

x=0.216837257 ?

OpenStudy (anonymous):

ohh so this is the exact form? 6x+8 = tan^-1(1/7)

OpenStudy (agent0smith):

See exact form above :P

OpenStudy (anonymous):

oh.. so this is the exact form? x = (tan^-1(1/7) - 8)/6 and this is the decimal form ? x=0.217 ?

OpenStudy (agent0smith):

x=-1.309683+kPi/6, k is the exact form, from @purger

OpenStudy (phi):

be sure to be in radian mode

OpenStudy (phi):

and round to 3 decimal places

OpenStudy (agent0smith):

forgot the kpi too x = (tan^-1(1/7) - 8 + k*pi)/6

OpenStudy (anonymous):

ohh okay :P was in the wrong mode :P so exact form: x = (tan^-1(1/7) - 8 + k*pi)/6 and decimal form: -1.310 ?? is that right now? :)

OpenStudy (agent0smith):

^decimal form needs kpi/6 too

OpenStudy (anonymous):

ohh :/ so decimal form= -1310kpi ____________ 6 ??

OpenStudy (anonymous):

is that how it's written? or is it -1.310(kπ/6) ? and oops i forgot the decimal point above.. haha it should be -1.310 :P

OpenStudy (agent0smith):

PLUS kpi/6 :P

OpenStudy (anonymous):

ohh ok :) yay! thank you!!! :)